OpenStudy (anonymous):

A and B are independent events. Find the missing probability. P(B)=1/5, P( A and B)=2/13. Find P(A)

OpenStudy (anonymous):

\[P(A)+P(B)=\frac{2}{13}\]\[P(A)+\frac{1}{5}=\frac{2}{13}\]I think it's like this.. If so solve to get P(A)..